Two workers have jobs as mowing lawns along the freeway. If one man takes three hours, and the second man takes six hours, how long will it take for both of the men to mow the lawn together?
Found 2 solutions by jorel555, ikleyn:
Answer by jorel555(1290)

Working together, the men can do:
1/3+1/6=3/6=1/2 the job per hour;
Thus it would take them 2 hours working together to finish 1 job. ☺☺☺☺

Answer by ikleyn(52814)